OpenStudy (anonymous):

suppose an isosceles triangle ABC has A=pi/4 and b=c=3. What is the length of a^2?

OpenStudy (owlcoffee):

Try using the pythagorean theorem.

OpenStudy (blockcolder):

Law of Cosines: \[a^2=b^2+c^2-2bc\cos{A}\]

OpenStudy (anonymous):

Answer is 3^2(2- sqrt2)